OpenStudy (anonymous):

Non-homogeneous recurrence solving. Need help with beginning. Question: tn = 2t(n-1) + n^2 The n of tn and 2t(n-1) is subscripted In my notes I need to find the characteristic equation first. All the examples say I need to start with the form tn - t(n-1) = 3^n <- always x^n on right side Here I made the equation into: tn - 2t(n-1) = n^2 n^2 isn't of exponent n and I'm unsure if this is a problem. The next steps include eliminating the right side by replacing the n with n-1 and multiplying the equation such that you can cancel the right side out by subtracting the original eq

OpenStudy (freckles):

\[t_n-c t_{n-1}=f(n) \\ u_{n}t_n-c u_{n}t_{n-1}=f(n) u_{n} \text{ multiply both sides by } u_{n} \\ \\ \text{ this is a sequence of my choice } \ \text{ Let } c u_n=u_{n-1} \\ \\ c=\frac{u_{n-1}}{u_n} \\ \text{ product of both sides } \\ c \cdot c \cdot c \cdots c=\frac{u_0}{u_1} \cdot \frac{u_1}{u_2} \cdot \frac{u_2}{u_3} \cdots \frac{u_{n-1}}{u_n} \\ \\ c^n=\frac{u_0}{u_n} \\ u_n=u_0 c^{-n}\] So let's back up a little... \[u_n t_n-u_{n-1}t_{t-1}= f(n)u_n \\ \\ \text{ sum both sides from } n=1 \text{ to } n=N \\ (u_1 t_1-u_0t_0)+(u_2 t_2-u_1 t_1)+(u_3t_3-u_2 t_1)+ \cdots +(u_N t_N-u_{N-1}t_{N-1}) \\ = \sum_{n=1}^{N} f(n) u_n \\ \implies -u_0 t_0+u_N t_N =\sum_{n=1}^{N} f(n) u_n \\ \implies u_N t_N= \sum_{n=1}^{N} f(n) u_n+u_0 t_0\] So pluggin u(n) we have and solving for t(N)... \[t_N= \frac{1}{u_{N}}( \sum_{n=1}^{N} f(n) u_n +u_0 t_0 )\\ t_N=\frac{c^{N}}{u_0} (\sum_{n=1}^{N} f(n) u_0 c^{-n}+u_0t_0) \\ t_N= c^{N} (\sum_{n=1}^{N} f(n) c^{-n} +t_0) \\ t_N= c^{N} \sum_{n=1}^N (f(n) c^{-n}) +c^N t_0\]

OpenStudy (freckles):

though you could also use undetermined coefficient way ... that above sum may be easy to simplify depending on f(n) could be hard though

OpenStudy (freckles):

not sure what you mean with n^2 thingy you wrote

OpenStudy (freckles):

I don't know what steps you are trying to describe either

OpenStudy (freckles):

example say f(n)=n and c=1 so we have \[t_N=1^{N} \sum_{n=1}^{N} (n 1^{-n})+1^{N} t_0 \\ t_N=\sum_{n=1}^{N} n+t_0 \\ t_N=\frac{N(N+1)}{2}+t_0 \\ \text{ replace }N \text{ with } n \\ t_n=\frac{n(n+1)}{2}+t_0 \text{ this is solution to } t_n-t_{n-1}=n\]

OpenStudy (freckles):

or you could instead of using that formula above use undetermined coefficients way.. \[t_n-t_{n-1}=n \\ \text{ first find homoegenous solution } t_n-t_{n-1}=0 \\ \text{ Let } t_n=Cr^n \\ \text{ so } t_{n-1}=Cr^{n-1} \\ Cr^n-Cr^{n-1}=0 \\ \text{ assume } r,C \neq 0 \\ \text{ divide both sides by } Cr^n \\ 1-r^{-1}=0 \implies 1=r^{-1} \implies 1=r \\ \text{ so homogeneous solution is } t_n=C (1)^n \text{ or just } t_n=C\] now time to guess the particular solution my guess is that it will be linear and I'm basing this off of the right hand expression which is n But we only need to really look at An instead of An+B since we already have a constant as a part of our solution (see homogeneous solution) \[t_n-t_{n-1}=n \\ \text{ Let } t_n=A n \\ t_{n-1}=A(n-1) \\ \text{ plug \in } \\ An-A(n-1)=n \\ \text{ distribute } \\ An-An+A=n \\ \text{ so } t_n=An \text{ fails } \\ \text{ \let's try } t_n=An^2+Bn \\ t_{n-1}=A(n-1)^2+B(n-1) \\ An^2+Bn-A(n-1)^2-B(n-1)=n \\ An^2+Bn-An^2+2An-A-Bn+B=0 \\ 2An+(-A+B)=n \\ \implies -A+B=0 ; 2A=1 \\ A=\frac{1}{2} \\ B=\frac{1}{2} \\ \\ \] so the final solution should include both homogeneous and particular solution \[t_n=\frac{1}{2}n^2+\frac{1}{2}n+C\]

OpenStudy (freckles):

you can also use generating function

OpenStudy (anonymous):

Thanks for all the information covering a bunch of different methods. The method I need to use appears to be: Find characteristic equation (homogeneous part): tn - 2t(n-1) = 0 n - 2 = 0 n = 2 Find general solution for non homogeneous part: Form of (n - b)^(d +1) where d is the degree of polynomial and b is from the part of the equation that I mentioned earlier that is said to have to be in the form x^n. So if it was 4^n, b = 4. However with n^2, I'm unsure what b value this takes or if I have to convert it to x^n somehow. Say b = 4 and d = 1 for this example, then (n - 4)^2 so the 2 parts combined = (n-2)(n-4)^2 so n = 2 n = 4 and n =4 So after finding the roots then I need to finish the question by solving for the constant coefficients.

OpenStudy (anonymous):

Roots would then be used as tn = c1(2^n) + c2(4^n) + c3(4^n)

OpenStudy (misssmartiez):

Classic sonic* c: @rebeccaxhawaii @mathmale @Directrix @ShadowLegendX @jim_thompson5910 1: I fear that I can't help with this one; sonnnniiikkkkku.... (>:D), however, those whom I named should help. 2: Isn't this chemistry?

OpenStudy (anonymous):

It's for an algorithms class. Thanks though!

OpenStudy (misssmartiez):

:) gotta go fast ;)

OpenStudy (freckles):

let me read your response real quick

OpenStudy (freckles):

so you only looking at one type of form for a particular solution is what I see from your post above

OpenStudy (freckles):

that form being: \[t_n=(n - b)^{d +1}\]

OpenStudy (anonymous):

@MissSmartiez You're too slow!

OpenStudy (anonymous):

tn=(n−b)d+1 + homogeneous part. Homogeneous part of that example was (n - 2)

OpenStudy (anonymous):

so homogeneous + non homogeneous

OpenStudy (freckles):

we need to compare this to the right hand side of the equation which is n^2 so that would mean we would want to choose \[t_n=(n-b)^{2+1}=(n-b)^3\] hmmm.... \[t_{n-1}=(n-1-b)^3=(n-b-1)^3=(n-b)^3-3(n-b)^2+3(n-b)-1 \\ \text{ so you have } t_n-2t_{n-1}=n^2 \\ \\ \text{ plug in } \\ (n-b)^3-2[(n-b)^3-3(n-b)^2+3(n-b)-1]=n^2 \\ -(n-b)^3+6(n-b)^2-6(n-b)+2=n^2 \\ \] this form doesn't seem to be working...

OpenStudy (freckles):

it looks like you are trying to do this by guessing the particular solution though so I might choose to go with \[t_n=An^2+Bn+C\]

OpenStudy (freckles):

since the right hand side is n^2

OpenStudy (freckles):

if the right hand side side was n^3 I would write a 3rd degree polynomial with 4 unknown constant values

OpenStudy (freckles):

http://web.cs.wpi.edu/~cs504/s00m/notes/recurrence/solve/step2/step2.html i love this website when I was learning what guess I should choose for the particular solution

OpenStudy (freckles):

hey @Its_Sonic I think I might have figured out the way you are trying to do

OpenStudy (freckles):

\[t_n-2 t_{n-1}=1^n n^2\] is a way you can express your equation this means b=1 and d=2

OpenStudy (freckles):

\[(n-2)(n-1)^3 =0 \text{ this is our new characteristic equation }\]

OpenStudy (freckles):

the (n-2) came from the homogeneous part and the (n-1)^3 came from the right hand side of the equation which will give us our particular solution

OpenStudy (freckles):

\[t_n=c_1 2^n+c_2 1^n+c_2 x 1^n+c_3 x^2 1^n \\ \text{ or we can write as } \\ t_n=c_1 2^n +c_2+c_2x+c_3 x^2\]

OpenStudy (freckles):

tell me what you think when you want

OpenStudy (anonymous):

Ahhh that actually makes sense to me now. The right side is in form b^n . p(n) but I kept seeing only b^n in examples without the polynomial part. Because it was missing entirely I didn't think n^2 was just the p(n) portion.

OpenStudy (freckles):

I never seen this way so sorry for slowness

OpenStudy (anonymous):

What strategy would be best to solve for the constants? Notes use Gaussian elimination for one.

OpenStudy (freckles):

I usually plug in the initial conditions and solve any resulting system of equations and yes Gaussian elimination is a method it is probably much better that the more elementary ways that I normally use :p so basically I don't have a better method of offer you :( but that doesn't mean one doesn't exist